OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). sin^2x + sin x = 0

sam (.sam.):

\[\sin^2x + \sin x = 0 \\ \\ (\sin(x))(\sin(x) + 1) = 0\] Then you'll say \[\sin(x)=0~~~~~~~~and~~~~~~~\sin(x)+1=0\] Solve 

OpenStudy (anonymous):

is it 0, pi, 3pi/2

sam (.sam.):

Correct

OpenStudy (anonymous):

thanks!

sam (.sam.):

yw